OpenStudy (anonymous):

find the slope of the line through the pair of points (6,12) and (-6,-2)

OpenStudy (austinl):

Ok, so the formula for slope is this. \[m=\frac{y_2-y_1}{x_2-x_1}\] I would use the first ordered pair as the y&x sub 2's. So using this, would you care to make an attempt?

OpenStudy (anonymous):

(0, 5)

OpenStudy (anonymous):

\[\frac{ -2+12 }{ -6+6}=\frac{ -14 }{ 6 }=\frac{ 7 }{ 6 }\]

OpenStudy (anonymous):

dawijr i got the same thing (0,5)

OpenStudy (austinl):

I get, \[m=\frac{12-(-2)}{6-(-6)}\] \[m=\frac{14}{12}\] Which simplified is \[m=\frac{7}{6}\] Very good!

OpenStudy (anonymous):

so (0,5) is right austinL

OpenStudy (austinl):

The problem presented asked for the slope of a line that went through the two points given. (0,5) is another set of coordinates.

OpenStudy (anonymous):

oooooooooohhh ok i get now

OpenStudy (anonymous):

0,5 is the midpoint

OpenStudy (anonymous):

yes

OpenStudy (austinl):

The problem didn't ask for midpoint, and that midpoint is under the assumption that the line is from (6,12) to (-6,-2). The problem says that the line passes through the points which would lead me to believe that the line extends infinitely in each direction.

OpenStudy (anonymous):

i said the same thing

OpenStudy (anonymous):

Ok well she got the answer so........